OpenStudy (ulisses23):

What is the y-intercept of the function f(x)=4 – 5x?

OpenStudy (kittiwitti1):

When a function is in the form of slope-intercept, \(y=mx+b\), the y-intercept will be the \(b\) value. Here your function has been arranged in \(y=b-mx\)... so you need to rearrange it. Hint:\[f(x)=4-5x\longleftrightarrow y=b-mx\]

OpenStudy (kittiwitti1):

Oops... I mean, \(b+mx\). The slope \(m\) can be negative.

OpenStudy (ulisses23):

thank you !!

OpenStudy (kittiwitti1):

You are very welcome :-) Good luck on your studies, ♣

OpenStudy (kittiwitti1):

Don't forget to close this problem once you are done with it, :-)

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)